# Aligning Speech to Languages to Enhance Code-switching Speech Recognition

### v2

@@ III-C Balancing training via toke-level language weights
While computing the number of speech frames for each language is challenging, the token distribution within a code-switching corpus can be assessed before training. Therefore, we propose to address the imbalance issue by incorporating language weights in (12), being a weighted cross-entropy function as
@@
where $w^{\mathrm{lang}}_{c}$ denotes the normalized language weight of language $c$ and can be made inversely proportional to its token count in the training data. However, the token ratio may not align precisely with speech frames due to variations in speech rates across languages. We therefore propose to tune them initially based on the token ratio and accounting for speech rate variations when applying language weights to the LAL. The relationship between the initial language weights can be illustrated by
@@
With the above, the CS-ASR model is optimized via an objective function being the sum of CTC loss, encode-decoder loss, and the proposed LAL such that
@@ III-D Linguistic hint for prompting LLM
With reference to Fig. 1, we propose to adopt the LLM-based generative error correction method to improve CS-ASR [29]. To fine-tune the LLM efficiently, LoRA is employed while the LLM is kept frozen during training as illustrated in Section II-B. The prompt originally designed in [29] is used as shown in Fig. 4(a), where the N-best hypotheses are extracted from the ASR output before being inserted in the prompt. The LLM is then optimized to deduce the correct transcript, which is the ground-truth transcription shown in Fig. 4 during training, by leveraging the information provided in these hypotheses. However, code-switching gives rise to more intricate token alternatives with similar auditory or syntactic characteristics compared to a monolingual application. This complexity persists when performing generative error correction for ASR hypotheses. Inspired by the use of chain-of-thought as additional supervision when fine-tuning LLM for downstream tasks [44, 28], we propose to employ an additional linguistic hint during prompting to address the aforementioned challenge in CS-ASR.
@@
The linguistic hint is then inserted in the used prompt during fine-tuning as shown in Fig. 4(b). We propose two types of hints—the monolingual hint, where only $<$language id$>$ words are included in the transcription, and the code-switching hint, where speech is multilingual and that $<$both language ids$>$ words are included in the transcription.
@@ IV-A Datasets
We conducted the experiments on data extracted from the ASRU 2019 Mandarin-English code-switching speech recognition challenge and the SEAME dataset [24, 25]. The ASRU 2019 Mandarin-English code-switching speech recognition challenge consists of four datasets, including a 500-hour Mandarin-only training set, a 200-hour intra-sentence English-Mandarin code-switching training set, a 40-hour intra-sentence English-Mandarin code-switching development set, and a 20-hour intra-sentence English-Mandarin code-switching test set. In the experiments, the models were trained on the 200-hour CS training set, validated on the development set, and evaluated on the test set. The SEAME dataset, on the other hand, is a Mandarin-English code-switching corpus containing spontaneous conversational speech [24]. This dataset encompasses both intra- and inter-sentence code-switching speech. We divided the SEAME dataset into a 96.6-hour training set, a 4.9-hour validation set, and two test sets being denoted as $\text{test}_{\texttt{man}}$ and $\text{test}_{\texttt{sge}}$ following the same partitioning method described in [3]. Detailed information about the test sets is provided in Table I, while Table II highlights the duration ratio of each language, where the language labels are annotated at the utterance level.
While both datasets involve English-Mandarin code-switching, the primary distinction between them lies in the accent. The ASRU dataset, recorded in mainland China, is characterized by a dominant Chinese accent and text. In the training and development sets, each sentence, on average, consists of 8.6 Chinese characters and 1.6 English words. In contrast, the SEAME dataset comprises audio recordings from Singapore and Malaysia, featuring South-East Asian accents. Additionally, code-switching occurs more frequently within the SEAME dataset compared to the ASRU data due to the bilingual education and language policies in Singapore and Malaysia [45]. This suggests that the SEAME data might pose greater challenges for CS-ASR compared to the ASRU data.
@@
TABLE II: Duration ratio for Mandarin, English, and code-switching speech of test sets in ASRU and SEAME data. The language annotations are at utterance level
Subset Duration ratio (%)
Mandarin English Code-Switching
ASRU test 0 0 100
$\text{SEAME test}_{\texttt{man}}$ 14 7 79
$\text{SEAME test}_{\texttt{sge}}$ 6 41 53
@@ IV-B Data preprocessing
Since the SEAME dataset contains a small training set of about 98 hrs, we augmented the training data using speed perturbation and SpecAugment [46, 47] for training models from scratch. Two training strategies were adopted, where one develops the model on data without speed perturbation while the other trains the model on the entire augmented data. The speech perturbation was applied with factors 0.9, 1.0, and 1.1. With respect to the ASRU dataset, only SpecAugment was applied for data augmentation. SpecAugment adopted the default setup in ESPnet for two datasets [2]. The time-warp mask size was set to five, and two time and frequency masks were applied, with their lengths uniformly selected from the range of 0 to 40 for time masks and 0 to 30 for frequency masks. Speech samples within both corpora were segmented into durations ranging from 0.1 to 20 s. We extracted $F=80$ dimensional log-Mel-Fbank features for each speech segment before applying the cepstral mean and variance normalization.
For the train-from-scratch models, we employed BPE to tokenize the English words in the two English-Mandarin code-switching corpora and split all Mandarin words into individual characters. For the SEAME dataset, this resulted in a total of $V=5,628$ tokens, including 3,000 English BPE tokens, 2,624 Mandarin characters, and four special tokens for $<$unk$>$, $<$noise$>$, $<$blank$>$, and $<$sos/eos$>$. For the ASRU data, the same tokenization process yielded a total of $V=6,923$ tokens, including 3,000 English BPE tokens, 3,920 Mandarin characters, and three special tokens for $<$unk$>$, $<$blank$>$, and $<$sos/eos$>$. The variable $C=3$ denotes three language classes converted from special, English, and Mandarin tokens.
We fine-tuned the Chinese LLaMA-2 on a subset of the SEAME training set comprising approximately 60,000 speech segments and the development set of the ASRU data comprising approximately 20,000 speech segments, respectively. N-best lists of the above speech segments were subsequently extracted and incorporated into the prompts. In addition, we removed the $<$noise$>$ and $<$unk$>$ labels from the N-best list since the LLM can hardly address these special tokens.
For experiments involving Whisper models, we adopted the data preprocessing and tokenization methods outlined in [37]. Force alignment is utilized to generate frame-level language annotations and is performed using the fine-tuned Whisper-small (FA-LB) and Whisper-large (FA-UB) models.
@@
IV-C Model configuration and experiment setup
The baseline model is a Conformer-based hybrid CTC/Attention ASR model comprising twelve Conformer encoder layers, six Transformer decoder layers, and a CTC module. The Conformer employs the macaron structure [31], where its convolutional neural network (CNN) module has a kernel size of 15. The swish activation function is applied [48]. A CNN layer first subsamples the input features and projects them into $D=256$ dimensions before feeding into the macaron modules. All attention layers within the encoder and decoder modules have four attention heads with input and output dimensions being $D=256$, and the inner layer of the position-wise feed-forward network is of dimensionality 2048.
The bi-encoder and language posterior bias methods have also been implemented as baselines [11, 4]. In the bi-encoder CS-ASR system, we replicated the encoder module of the baseline model described in the previous paragraph to form the bi-encoder module. The outputs of these two encoders are fed into a linear layer, which serves as a built-in LID module with an output dimensionality of two (corresponding to two languages), before calculating the weighted encoder outputs. It is useful to note that the encoder modules of the bi-encoder method are pre-trained on English and Mandarin corpora separately. Nevertheless, this pre-training was deliberately omitted to ensure a fair comparison. For the language posterior bias-based (LPB) CS-ASR model, the model configuration recommended in [4] was used. The LPB system closely resembles the baseline hybrid CTC/attention model, with an additional language diarization decoder. This language diarization decoder mirrors the structure of the ASR decoder.
@@ IV-C Model configuration and experiment setup
Experiments associated with the proposed linguistic hint utilized a Chinese LLaMA2-7B model  [^1] https://huggingface.co/hfl/chinese-llama-2-7b. The 5-best hypotheses were first extracted from the ASR output before being used in the prompt to fine-tune the LLM. Fine-tuning was performed via LoRA with rank $r=4$ and the AdamW optimizer for ten epochs with a batch size of 128 on the SEAME data, where LoRA is applied to the query and value modules within the self-attention modules. The learning rate was increased from 0 to 0.0002 over 100 update steps, followed by a linear decay. During inference, a temperature of 0.7 was applied to allow for the creativity of the LLM, while other hyper-parameters were fixed at their default settings. We evaluated all methods by employing the MER comprising WER for English and CER for Mandarin.
Experiments relevant to Whisper models utilized the Whisper-small model [37]. Fine-tuning was conducted using an AdamW optimizer over five epochs, with a learning rate increased from 0 to $1\times 10^{-5}$ over 10,000 update steps, followed by a linear decay. The training employed a batch size of eight samples, with gradients accumulated every two updates. Inference was carried out using the best-performing model on the development sets, with a beam size of one. Language prompts $<$zh$>$ and $<$en$>$ are used for the ASRU and SEAME data, respectively. Since Whisper models were pre-trained on large-scale data, $\beta$ is set to 0.05 and 0.01 for ASRU and SEAME data, respectively, to prevent the LAL or frame-level LID from exerting excessive influence during fine-tuning. The frame-to-token alignment for LAL is computed via the cross-attention matrices within the last decoder layer. LoRA was applied to the query, key, value, and linear modules with rank $r=8$ during parameter-efficient fine-tuning.
@@ V-A Impact of $\beta$ values
We next conduct experiments on the SEAME dataset without using speed perturbation during training and results are presented in Table IV. Similar to the ASRU dataset, the hybrid CTC/attention model that incorporates the proposed LAL outperforms the vanilla configuration. In addition, the hybrid CTC/attention model with LAL exhibits consistently high performance for $1.0\leq\beta\leq 3.0$. This robustness on various datasets and $\beta$ values indicates the effectiveness of the proposed LAL. As opposed to results achieved on the ASRU dataset, the highest overall performance on the SEAME data is achieved when $\beta=0.1,0.5,$ and 3.0. Due to the different duration ratios of languages of the two test sets, the above implies that a lower $\beta$ value leads to a higher performance on the dataset containing predominantly monolingual data.
The impact of $\beta$ values can also be discussed from the perspective of accent. Compared to the ASRU data, which is characterized by a dominant Chinese accent, the SEAME data is characterized by Southeast Asian accents. Due to the impact of the Southeast Asian accents on the two languages, the English and Mandarin speech in the SEAME dataset can exhibit auditory similarity, making LID for the SEAME data more challenging than for the ASRU data. Therefore, a larger $\beta$ value is required to achieve high performance when training the CS-ASR model on the SEAME data.
@@
TABLE V: Performance of the proposed method with $\beta=1.5$ and different language weights to balance the Mandarin-dominant ASRU data during training in terms of English WER, Mandarin CER, and MER (%). “$ft$” denotes fine-tuning
@@ V-B Balancing the ASRU dataset during training
The effect of language weights employed in (13) for balancing the Mandarin-dominant ASRU data is shown in Table V. It is worth noting that, similar to the number of English frames, the number of frames for the class “other” is also significantly lower than that for the Mandarin frames. Since the class “other” does not contribute to the language identities in the CS-ASR task, this class is not balanced during training and the weight is always set to $w_{o}ther^{\mathrm{lang}}=1$ except for the learnable language weight. In this work, $w_{c}^{\mathrm{lang}}=1$ for all classes is the default setup.
@@ V-C Impact of speed perturbation
As discussed in Section V-B, the frame-level language identification can be affected by speech rate. We, therefore, investigate the impact of the speed perturbation on the CS-ASR performance when LAL is incorporated. The experiments were conducted on the SEAME dataset and in contrast to the experiments shown in Table IV, we utilize speed perturbation to augment the training data with results being summarized in Table VI.
Results presented show that the hybrid CTC/attention model achieves higher performance after employing speed perturbation as data augmentation. However, due to speech rate variations, one token can be aligned with a different number of acoustic frames when speed perturbation is applied. Therefore, further incorporation of the proposed LAL may lead to lower performance. In addition, the hybrid CTC/attention model with LAL being incorporated achieves the highest performance with $\beta=3$. This is consistent with the performance of its counterpart on the SEAME dataset without speed perturbation. It is useful to note that the proposed method achieves the highest performance on the ASRU dataset when $\beta=1.5$. Since a higher $\beta$ value is required for the SEAME data to achieve the highest performance, this also underpins that the SEAME dataset presents greater challenges in language discrimination than the ASRU dataset.
In addition, when speech perturbation is not used, the results presented in Tables III and IV show that the proposed LAL can consistently improve the CS-ASR performance compared to the baseline system. Similar to other hyper-parameters such as learning rate and $\alpha$ in (7) and (15) which require fine-tuning to achieve optimal performance, a fine-tuned $\beta$ results in significant improvements over the baseline model.
TABLE VII: Performance evaluation of the proposed method and state-of-the-art approaches on test data from the ASRU 2019 challenge and SEAME dataset in terms of English WER, Mandarin CER, and MER (%). “PT” denotes the model is pre-trained, “$ft$” denotes fine-tuning, frame-level labels for “LID” are generated via force alignment, “FA-LB” and “FA-UB” denote the force alignment computed via fine-tuned Whisper-small (lower bound) and Whisper-large (upper bound) models, respectively

### v3

@@ III-C Balancing training via token-level language weights
While computing the number of speech frames for each language is challenging, the token distribution within a code-switching corpus can be assessed before training. Therefore, we propose to address the imbalance issue by incorporating language weights in (12), resulting in a weighted cross-entropy function
@@
where $w^{\mathrm{lang}}_{c}$ denotes the normalized language weight of language $c$ and is inversely proportional to its token count within the training data. However, the token ratio may not align precisely with speech frames due to variations in speech rates across languages. We therefore propose to tune them initially based on the token ratio and accounting for speech rate variations when applying language weights to the LAL. The relationship between the initial language weights is given by
@@
With the above, the CS-ASR model is optimized via an objective function being the sum of CTC loss, encoder-decoder loss, and the proposed LAL such that
@@ III-D Linguistic hint for prompting LLM
With reference to Fig. 1, we propose to adopt the LLM-based generative error correction method to improve CS-ASR [31]. To fine-tune the LLM efficiently, LoRA is employed while the LLM is kept frozen during training, as illustrated in Section II-B. The prompt originally designed in [31] is used as shown in Fig. 4(a), where the N-best hypotheses are extracted from the ASR output before being inserted in the prompt. The LLM is then optimized to deduce the correct transcript, which is the ground-truth transcription shown in Fig. 4 during training, by leveraging the information provided in these hypotheses. However, code-switching gives rise to more intricate token alternatives with similar auditory or syntactic characteristics compared to a monolingual application. This complexity persists when performing generative error correction for ASR hypotheses. Inspired by the use of chain-of-thought as additional supervision when fine-tuning LLM for downstream tasks [46, 30], we propose to employ an additional linguistic hint during prompting to address the aforementioned challenge in CS-ASR.
@@
The linguistic hint is then inserted into the used prompt during fine-tuning as shown in Fig. 4(b). We propose two types of hints—the monolingual hint, where only <<language id>> words are included in the transcription, and the code-switching hint, where speech is multilingual and that <<both language ids>> words are included in the transcription.
@@ IV-A Datasets
We conducted the experiments on data extracted from the ASRU 2019 Mandarin-English code-switching speech recognition challenge and the SEAME dataset [26, 27]. The ASRU 2019 Mandarin-English code-switching speech recognition challenge consists of four datasets, including a 500-hour Mandarin-only training set, a 200-hour intra-sentence English-Mandarin code-switching training set, a 40-hour intra-sentence English-Mandarin code-switching development set, and a 20-hour intra-sentence English-Mandarin code-switching test set. In the experiments, the models were trained on the 200-hour CS training set, validated on the development set, and evaluated on the test set. The SEAME dataset, on the other hand, is a Mandarin-English code-switching corpus containing spontaneous conversational speech [26]. This dataset encompasses both intra- and inter-sentence code-switching speech. We divided the SEAME dataset into a 96.6-hour training set, a 4.9-hour validation set, and two test sets denoted by $\text{test}_{\texttt{man}}$ and $\text{test}_{\texttt{sge}}$ following the same partitioning method described in [4]. Detailed information regarding the test sets is provided in Table I. Table II highlights the duration ratio of each language, where the language labels are annotated at the utterance level. Example text sentences from the two datasets are shown in Table XII.
While both datasets involve English-Mandarin code-switching, the primary distinction between them lies in the accent. The ASRU dataset, recorded in mainland China, is characterized by a dominant Chinese accent and text. In the training and development sets, each sentence, on average, consists of 8.6 Chinese characters and 1.6 English words. In contrast, the SEAME dataset comprises audio recordings from Singapore and Malaysia, featuring South-East Asian accents. On average, sentences in the SEAME training and development sets contain 9.5 Chinese characters and 4.4 English words. Additionally, code-switching occurs more frequently within the SEAME dataset compared to the ASRU data due to the bilingual education and language policies in Singapore and Malaysia [47]. This suggests that the SEAME data might pose greater challenges for CS-ASR compared to the ASRU data.
@@
TABLE II: Utterance-level duration ratios and dataset-level token distribution (English words and Mandarin characters) ratios in the ASRU and SEAME test sets
Subset Duration ratio (%) Token ratio (%)
Man Eng CS Man Eng
ASRU test 0 0 100 89 11
$\text{SEAME test}_{\texttt{man}}$ 14 7 79 74 26
$\text{SEAME test}_{\texttt{sge}}$ 6 41 53 37 63
@@ IV-B Data preprocessing
Since the SEAME dataset contains a small training set of approximately 98 hrs, we augmented the training data using speed perturbation and SpecAugment [48, 49] for training models from scratch. Two training strategies were adopted, wherein one develops the model on data without speed perturbation while the other trains the model on the entire augmented data. The speech perturbation was applied with factors 0.9, 1.0, and 1.1. With respect to the ASRU dataset, only SpecAugment was applied for data augmentation. SpecAugment adopted the default setup in ESPnet for two datasets [2]. The time-warp mask size was set to five, and two time and frequency masks were applied, with their lengths uniformly selected from the range of 0 to 40 for time masks and 0 to 30 for frequency masks. Speech samples within both corpora were segmented into durations ranging from 0.1 to 20 s. We extracted $F=80$ dimensional log-Mel-Fbank features for each speech segment before applying the cepstral mean and variance normalization.
For the train-from-scratch models, we employed BPE to tokenize the English words in the two English-Mandarin code-switching corpora and split all Mandarin words into individual characters. For the SEAME dataset, this resulted in a total of $V=5,628$ tokens, including 3,000 English BPE tokens, 2,624 Mandarin characters, and four special tokens (<<unk>>, <<noise>>, <<blank>>, and <<sos/eos>>). For the ASRU data, the same tokenization process yielded a total of $V=6,923$ tokens comprising 3,000 English BPE tokens, 3,920 Mandarin characters, and three special tokens (<<unk>>, <<blank>>, and <<sos/eos>>). The variable $C=3$ denotes three language classes converted from special, English, and Mandarin tokens.
We fine-tuned the Chinese LLaMA-2 on a subset of the SEAME training set comprising approximately 60,000 speech segments and the development set of the ASRU data comprising approximately 20,000 speech segments. N-best lists of the above speech segments were subsequently extracted and incorporated into the prompts. In addition, we removed the <<noise>> and <<unk>> labels from the N-best list since the LLM can hardly address these special tokens.
For experiments involving Whisper models, we adopted the data preprocessing and tokenization methods outlined in [39]. Forced alignment is utilized to generate frame-level language annotations and is performed using the fine-tuned Whisper-small (FA-LB) and Whisper-large (FA-UB) models.
@@
IV-C Model configuration and experimental setup
The baseline model is a Conformer-based hybrid CTC/Attention ASR model comprising twelve Conformer encoder layers, six Transformer decoder layers, and a CTC module. The Conformer employs the macaron structure [33], where its convolutional neural network (CNN) module has a kernel size of 15. The swish activation function is applied [50]. A CNN layer first subsamples the input features and projects them into $D=256$ dimensions before feeding into the macaron modules. All attention layers within the encoder and decoder modules have four attention heads with input and output dimensions being $D=256$, and the inner layer of the position-wise feed-forward network is of dimension 2048.
The bi-encoder and language posterior bias methods have also been implemented as baselines [13, 5]. In the bi-encoder CS-ASR system, we replicated the encoder module of the baseline model described in the previous paragraph to form the bi-encoder module. The outputs of these two encoders are fed into a linear layer, which serves as a built-in LID module with an output dimension of two (corresponding to two languages), before calculating the weighted encoder outputs. It is useful to note that the encoder modules of the bi-encoder method are pre-trained on English and Mandarin corpora separately. Nevertheless, this pre-training was deliberately omitted to ensure a fair comparison. For the language posterior bias-based (LPB) CS-ASR model, the model configuration recommended in [5] was used. The LPB system closely resembles the baseline hybrid CTC/attention model, with an additional language diarization decoder. This language diarization decoder mirrors the structure of the ASR decoder.
@@ IV-C Model configuration and experimental setup
Experiments associated with the proposed linguistic hint utilized a Chinese LLaMA-2-7B model [^1] https://huggingface.co/hfl/chinese-llama-2-7b. The 5-best hypotheses were first extracted from the ASR output before being used in the prompt to fine-tune the LLM. Fine-tuning was performed via LoRA with rank $r=4$ and the AdamW optimizer for ten epochs with a batch size of 128 on the SEAME data, where LoRA is applied to the query and value modules within the self-attention modules. The learning rate was increased from 0 to 0.0002 over 100 update steps, followed by a linear decay. During inference, a temperature of 0.7 was applied to allow for the creativity of the LLM, while other hyper-parameters were fixed at their default settings.
Experiments relevant to Whisper models utilized the Whisper-small model [39]. Fine-tuning was conducted using an AdamW optimizer over five epochs, with a learning rate increased from 0 to $1\times 10^{-5}$ over 10,000 update steps, followed by a linear decay. The training employed a batch size of eight samples, with gradients accumulated every two updates. Inference was carried out using the best-performing model on the development sets, with a beam size of one. Language prompts <<zh>> and <<en>> were used for the ASRU and SEAME data, respectively. Since Whisper models were pre-trained on large-scale data, $\beta$ was set to 0.05 and 0.01 for ASRU and SEAME data, respectively, to prevent the LAL or frame-level LID from exerting excessive influence during fine-tuning. The frame-to-token alignment for LAL is computed via the cross-attention matrices within the last decoder layer. LoRA was applied to the query, key, value, and linear modules with rank $r=8$ during parameter-efficient fine-tuning.
@@ V-A Impact of $\beta$ values
We next conduct experiments on the SEAME dataset without using speed perturbation during training and present the results in Table IV. Similar to the ASRU dataset, the hybrid CTC/attention model that incorporates the proposed LAL outperforms the vanilla configuration. In addition, the hybrid CTC/attention model with LAL exhibits consistently high performance for $1.0\leq\beta\leq 3.0$. This robustness on various datasets and $\beta$ values indicates the effectiveness of the proposed LAL. As opposed to results achieved on the ASRU dataset, the highest overall performance on the SEAME data is achieved when $\beta=0.1,0.5,$ and 3.0. Due to the different duration ratios of languages of the two test sets, the above implies that a lower $\beta$ value leads to a higher performance on the dataset containing predominantly monolingual data.
The impact of $\beta$ values can also be discussed from the perspective of accent. Compared to the ASRU data, which is characterized by a dominant Chinese accent, the SEAME data is characterized by Southeast Asian accents. Due to the impact of Southeast Asian accents on the two languages, the English and Mandarin speech within the SEAME dataset can exhibit auditory similarity, making LID for the SEAME data more challenging than that for the ASRU data. Therefore, a larger $\beta$ value is required to achieve high performance when training the CS-ASR model on the SEAME data.
@@
TABLE V: Performance of the proposed method with $\beta=1.5$ and different language weights to balance the Mandarin-dominant ASRU data during training in terms of English WER, Mandarin CER, and MER (%). “ftft” denotes fine-tuning
@@ V-B Balancing the ASRU dataset during training
The effect of language weights employed in (13) for balancing the Mandarin-dominant ASRU data is shown in Table V. We note that, similar to the number of English frames, the number of frames for the class “other” is also significantly lower than that for the Mandarin frames. Since the class “other” does not contribute to the language identities in the CS-ASR task, this class is not balanced during training and the weight is always set to $w_{other}^{\mathrm{lang}}=1$ except for the learnable language weight. In this work, $w_{c}^{\mathrm{lang}}=1$ is the default setup for all classes.
@@ V-C Impact of speed perturbation
As discussed in Section V-B, the frame-level language identification can be affected by speech rate. We, therefore, investigate the impact of the speed perturbation on the CS-ASR performance when LAL is incorporated. The experiments were conducted on the SEAME dataset and in contrast to the experiments shown in Table IV, we utilize speed perturbation to augment the training data with results summarized in Table VI.
These results show that the hybrid CTC/attention model achieves higher performance after employing speed perturbation as data augmentation. The proposed method achieves the highest performance on the ASRU dataset when $\beta=1.5$. Since a higher $\beta$ value is required for the SEAME data to achieve the highest performance, this also underpins that the SEAME dataset presents greater challenges in language discrimination than the ASRU dataset. While speed perturbation is effective for data augmentation and overall ASR performance improvement, it introduces inconsistencies in the alignment between text tokens and acoustic frames, where a single token may correspond to different numbers of frames under varying speech rates. These inconsistencies can reduce the reliability of frame-level supervision and the effectiveness of the proposed LAL. Therefore, further incorporation of the proposed LAL may lead to lower performance. Notwithstanding the above, LAL can still yield moderate performance improvement under speed perturbation when appropriately tuned, where the hybrid CTC/attention model with LAL being incorporated achieves the highest performance when $\beta=3$. This is consistent with the performance of its counterpart on the SEAME dataset without speed perturbation. Therefore, although the overall improvement may be less significant compared to training without perturbation, LAL remains beneficial and applicable in real-world CS-ASR scenarios, particularly when speed perturbation is not essential due to the availability of rich training data.
Furthermore, when speech perturbation is not used, the results presented in Tables III and IV show that the proposed LAL can consistently improve the CS-ASR performance compared to the baseline system. Similar to other hyper-parameters such as learning rate and $\alpha$ in (7) and (15) which require fine-tuning to achieve optimal performance, a fine-tuned $\beta$ results in significant improvements over the baseline model.
TABLE VII: Performance evaluation of the proposed method and state-of-the-art approaches on test data from the ASRU 2019 challenge and SEAME dataset in terms of English WER, Mandarin CER, and MER (%). “PT” denotes the model is pre-trained, “ftft” denotes fine-tuning, frame-level labels for “LID” are generated via forced alignment, “FA-LB” and “FA-UB” denote the forced alignment computed via fine-tuned Whisper-small (lower bound) and Whisper-large (upper bound) models, respectively